### v1

# Optimizing Charge Transport Simulation for Deep Learning Enhanced Spatial Resolution of the MÖNCH Detector

Abstract
Deep learning has the potential to significantly enhance the spatial resolution of hybrid detectors for electrons, as demonstrated for electron microscopy using 200 keV electrons with the MÖNCH 25 µm pitch charge integration pixel detector. The deep learning models have been trained using both simulation data, which are more straightforward to generate with varied parameters, and measurement data, which are more cumbersome to acquire due to the required specialized setup of the electron microscope. However, challenges arise when comparing simulation-based deep learning models to measurement-based models for electrons, as the spatial resolution achieved through simulations is notably inferior to that from measurements. Discrepancies are also observed when directly comparing X-ray simulations with measurements, particularly in the spectral output of single pixels. These observations collectively suggest that current simulations require optimization.
To address this, a Monte Carlo simulation was conducted to model the dynamics of charge carriers within the silicon sensor, aiming to refine charge transport modeling. The simulation encompasses the initial generation of the charge cloud, charge cloud drift, charge diffusion and repulsion, and electronic noise. The simulation results were validated with measurements from the MÖNCH detector for X-rays, and consistencies between the simulation and the measurements were significantly improved by accounting for the charge repulsion.
1 Introduction
The MÖNCH detector, currently under development at the Paul Scherrer Institut, is a charge-integrating hybrid pixel detector [1]. Its 25 µm pitch, combined with high dynamic range, fast readout, and low noise, demonstrates the potential of hybrid pixel detectors for more electron microscopy applications. Our prior study successfully demonstrated that deep learning can effectively reconstruct the incident position of single electron, thereby enhancing the spatial resolution for 200 keV electrons [2], where performance is primarily constrained by multiple scattering effects. Our deep learning models were trained using both simulation data and measurement data. Simulation data are more straightforward to generate with varied parameters, whereas measurement data acquisition are challenging due to the specialized setup required for the electron microscope. Additionally, simulations offer more detailed information, such as the electron track, which is not available in the measurements.
However, the small pitch of the MÖNCH detector not only enables high spatial resolution but also presents challenges in accurately modeling charge transport. These challenges become apparent when comparing simulation-based deep learning models to those based on measurements. The spatial resolution achieved using simulations is notably inferior to that achieved with measurements. To address the origin of this discrepancy, we conducted comparisons between measurements and simulations generated by [3] using the ProjectionPropagation module for 200 keV electrons and X-rays phohtons.
Figure 1 (a), (b), and (c) show the pixel energy spectra for 200 keV electron clusters and the pixel spectra of 3x3 clusters of copper and silver fluorescence, respectively. At higher energies, the simulation shows a larger number of entries than the measurement, as illustrated in the spectra and amplified in the ratio plots. This discrepancy suggests an underestimation of the charge-sharing effect in the simulation, which requires optimization. Furthermore, when comparing the pixel energy spectra for copper and silver X-rays (Figure 1(b) and (c)), the discrepancies are more pronounced for higher energy depositions.
This study aims to optimize the charge transport simulation for the MÖNCH detector. A time-stepping Monte Carlo simulation is conducted following the charge carrier dynamics in the silicon sensor. The simulation process includes the initial generation of the charge cloud by X-rays or electrons, charge cloud drift, charge diffusion and repulsion, and electronic noise. The charge repulsion, which is currently not considered in established simulation frameworks [3, 4, 5], is proved to be crucial for accurately reproducing the MÖNCH detector measurements. The following sections elaborate on the charge carrier dynamics, simulation algorithm, validation results, discussion, and future outlook.
Charge carrier dynamics
The dynamics of charge carriers in silicon sensors is governed by the continuity equation:
where represents the charge carrier density, is the diffusion coefficient, denotes the charge carrier mobility, and is the electric field. The generation and recombination of charge carriers are ignored in the depleted sensor volume. Driven by the bias voltage, the charge cloud, consisting of charge carriers, drifts towards the readout electrode. By neglecting charge repulsion, Equation 1.1 simplifies to , with the coordinate system’s origin at the center of the charge cloud. The solution to this equation is a Gaussian distribution with a standard deviation of in one dimension, a charge diffusion formula widely implemented in simulation frameworks.
However, the neglect of charge repulsion in the simulation leads to an underestimation of the charge cloud size, accounting for the discrepancies between the simulation and the measurements in Figure 1. Thus, charge repulsion must be taken into account in the simulation for more accurate modeling of charge transport in the MÖNCH detector. Given that mobility is a complex function of the electric field, finding analytical solutions to the continuity equation that account for charge repulsion is impractical. Therefore, this study developed a time-stepping Monte Carlo simulation to model charge carrier transport in the silicon sensor while incorporating charge repulsion. X-ray photon are simulated in this study due to their point-like energy deposition, which provides an ideal initial condition for simulations. Additionally, the accessibility in measurements using X-ray tubes facilitates direct comparisons.
2 Time-stepping Monte Carlo simulation
Boundry conditions
The absorption depth of X-rays in the silicon sensor follows an exponential distribution, with the attenuation length determined by the X-ray energy and the sensor material [10]. The absorbed energy is converted into charge at a conversion efficiency of 3.6 eV per electron-hole pair. The number of charge carriers is manually set to 100,000, evenly distributing the total charge. This is because the charge repulsion is the primary focus of this study, and a small number of charge carriers and a sparse charge cloud are sufficient apply the spherical symmetry assumption to calculate the electric field due to the charge repulsion.
The initial charge density distribution adheres to an empirical Gaussian distribution with µm [11], where is the energy deposition. The transport begins at the X-ray absorption depth and ends when the charge cloud center reaches the sensor pixel side.
Charge carrier mobility
Holes are collected in the MÖNCH detector. The widely-used Jacoboni-Canali model, which takes into account the saturation of the mobility for high electric fields, is employed to describe the mobility of holes in silicon as a function of the electric field and temperature [6]:
with phenomenological parameters for holes:
In this study, the temperature is set to 293 K, which is the cooling water temperature for the MÖNCH detector during measurements. The input electric field is determined by
where results from the bias voltage and arises from the charge repulsion, detailed in the following sections. The calculation assumes orthogonality between and , which is valid for charge carriers on the plane across the cloud center and is applied to the entire charge cloud. This assumption maintains spherical symmetry of the charge cloud and simplifies the calculation.
Drift of charge cloud
In a Cartesian coordinate system with the origin at the charge cloud center and the z-axis aligned with the drift direction, the electric field due to the bias voltage is expressed as:
where is the bias voltage, is the depletion voltage, and is the sensor thickness. The depletion voltage of the MÖNCH sensor is measured to be 30 V, and the sensor thickness is 320 µm. The applied bias voltage during measurements is 90 V.
The coordinate of the charge cloud center is updated as follows:
where is the average hole mobility of charge carriers, and is the time step, set to 0.01 ns.
Diffusion and repulsion of charge carriers
The diffusion and repulsion of charge carriers are considered in a Cartesian coordinate system with the origin at the center of the charge cloud. For a charge carrier at , diffusion in the direction is modeled as a random walk process:
where is the diffusion coefficient, and the sign indicates that the direction of diffusion is random. Similar diffusion processes are applied to the and directions.
The electric field due to the charge repulsion is calculated as:
where is the charge within the sphere of radius , is the vacuum permittivity, and is the relative permittivity of silicon. The assumption of spherical symmetry of the charge cloud simplifies the calculation of the electric field due to the charge repulsion. The coordinate update due to the repulsion, similar in the and directions, is given by:
Intermediate results and parameterization
Intermediate results from a simulated copper K X-ray with an absorption position of (0, 0, 2.5 µm) are presented below. Figure 1(a) plots the root mean square (RMS) of charge carrier coordinates over drift time with and without the charge repulsion. The RMS at the end of the simulation including the charge repulsion is 8.46 µm, which is significantly higher than the RMS without the repulsion (7.68 µm).
Parameterization of the charge distribution is crucial for efficiently generating simulation samples. The final charge distribution, with the repulsion considered, is fitted by a generalized Gaussian distribution:
where is the scale parameter, is the shape parameter, and is the gamma function. The fitted generalized Gaussian distribution is shown as the black curve in Figure 1(b). The shape parameter describes the broadening of the charge distribution due to charge repulsion. The of the fit is 1.0, indicating that the generalized Gaussian distribution effectively models the charge distribution. Good fits are also observed for different absorption depths and X-ray energies.
3 Validation of the simulation
The validation of the simulation is conducted by comparing the simulation results with measurements obtained using a MÖNCH detector for copper and silver X-ray fluorescence, specifically in terms of the pixel energy spectrum and the charge-weighted center of 3x3-pixel clusters. X-ray clusters of both the measurements and the simulation adher to the same selection criteria and processing steps.
Forming 3x3-pixel clusters
The simulation previously described is repeated for various absorption depths every 5 µm. One million X-rays are simulated, with the absorption depth sampled from an exponential distribution corresponding to the respective attenuation length and a random incident position uniformly distributed in the plane. For each X-ray, the charge carrier distribution relative to the cloud center is sampled from the fitted generalized Gaussian distribution for the nearest simulated absorption depth. Pixel noise, approximately as determined from measurements, is added to each pixel. The 3x3-pixel clusters are centered on the pixel with the maximum charge. Additionally, K X-rays are simulated following the same procedure, with the intensity ratio to K X-rays derived from an empirical study [7] for copper. In the case of the silver K the ratio was obtained from the double-Gaussian fit to K and K measurement.
Measurements were conducted using the MÖNCH detector with an X-ray tube featuring copper and silver targets. To process the raw data from the charge-integration MÖNCH detector, dark current subtraction was applied. Pixel-wise gain calibration was performed to convert the pixel readout from Arbitrary Digital Units to energy. The 3x3-pixel clusters were centered on the pixel with the maximum energy. Pile-up clusters were excluded from the analysis by ensuring no neighboring pixels of the cluster belonged to another cluster.
An 1 keV cluster energy selection around the K peak was applied to both data sets, further excluding noise and pile-up clusters.
Pixel energy spectrum
The pixel energy spectra of the MÖNCH detector for selected copper and silver X-ray clusters are illustrated in Figure 3 (a) and (b), respectively. Both figures show good agreement between the simulation and the measurements, with some remaining discrepancies. The simulation results exhibit a higher number of entries in lower energy bins, while the measurements show a higher number of entries in higher energy bins especially at the K peak. These discrepancies are primarily attributed to noise modeling, the non-linearity of the charge amplification, and cross-talk between the pixels, which causes a bump in the K peak visible only in the measurements. Overall, significant improvements are observed in the simulation results compared to results obtained without accounting for charge repulsion by comparing Figure 3 and Figure 1.
Charge weighted center ,
The charge weighted centers and are widely used in position interpolation to improve spatial resolution [8, 9]. For the 3x3-pixel clusters, the charge weighted centers and are calculated as follows:
Here denotes the energy of the pixel111Charge and energy are translated using a conversion coefficient of 3.6 eV per electron-hole pair. at the -th row and the -th column. With the origin of the coordinate system defined as the lower left corner of the center pixel, the charge weighted centers and range from 0 to 1.
The 2D histograms of the charge weighted centers and for copper and silver X-rays are depicted in Figures 4(a) and (b,) respectively. The z-axis in logarithmic scale represents the number of entries. The spectra after projection, obtained from both measurements and simulations, are displayed in Figure 4(c) and (d). The simulation aligns well with the measurements for copper X-rays, while there remain some discrepancies for silver X-rays. The sources of these discrepancies will be detailed in the following section.
4 Discussion, conclusion and outlook
The 25 µm pitch of the MÖNCH detector presents the challenge of a more accurate modeling of charge transport, while simultaneously offering a unique platform for the investigation. Meanwhile, we should note that the difference in the charge distribution with and without repulsion, as exampled in Figure 1(a), is less noticeable for hybrid pixel detectors with a larger pixel pitch, such as 55 µm of the Medipix detectors [12] and 75 µm of the Jungfrau detectors [13].
Sources of remaining discrepancies
While incorporating the charge repulsion into the simulation has significantly improved the agreement between the simulation and the measurements, some discrepancies persist. The simulation assumes a spherically symmetric charge distribution to simplify the calculation of the electric field due to charge repulsion, an assumption which does not strictly hold. Additionally, the gradient of the drift electric field is neglected in the simulation. When the charge cloud approaches the readout electrode, the symmetry of the charge cloud is broken. Furthermore, up to a 10% deviation in the mobility model exists [6].
In the measurements, uncertainties from measured parameters, such as the depletion voltage, contribute to these discrepancies. Non-linearity of charge amplification and cross-talk between pixels should also be considered, as discussed previously.
Conclusion and outlook
In this study, a time-stepping Monte Carlo simulation incorporating charge repulsion was developed to simulate charge carrier transport in the silicon sensor. Parameterization was implemented to effectively model the charge distribution and efficiently generate simulation samples. The simulation results were validated against measurements obtained from the MÖNCH detector for X-rays. By considering charge repulsion, the consistency between the simulation and the measurements has been significantly improved.
Further research is necessary to implement this approach for electron simulations. Modeling the charge cloud for electrons is more challenging than for X-rays due to the continuous energy deposition spectrum. The parameters of the charge cloud distribution should be modeled as a function of both the deposition depth and the deposition energy. Additionally, the repulsion between charge clouds must be taken into account. With optimized simulations, higher-quality simulation samples can be generated for deep learning applications. Various design parameters of the MÖNCH detector can be investigated to exploit the charge sharing effect.